OpenStudy (anonymous):

A new bike costs $250.00 and is marked 22% off, what is the sale price?

OpenStudy (jackellyn):

250*(22/100)=250*0.22=sale price

OpenStudy (anonymous):

i got 55

OpenStudy (jackellyn):

sorry its 250*0.22=how much deducted 250-55=sale price